The accompanying data are the number of wins and the earned run averages​ (mean number of earned runs allowed per nine innings​ pitched) for

eight

baseball pitchers in a recent season. Find the equation of the regression line. Then construct a scatter plot of the data and draw the regression line. Then use the regression equation to predict the value of y for each of the given​ x-values, if meaningful. If the​ x-value is not meaningful to predict the value of​ y, explain why not.

a) x=5 wins

b) x=10wins

c) x=21wins

d) x=15 wins

the equation of the regression line is y=___x+_____

round to two decimal places if necessary

construct a scatter plot of the data and draw the regression line.

​(a) Predict the ERA for

five

​wins, if it is meaningful. Select the correct choice below​ and, if​ necessary, fill in the answer box within your choice.

A) y=___ round to 2 decimal places as needed

B) it is not meaningful to predict this value of y because x=5 is not an x-value in the original data

C. it is not meaningful to predict this value of y because x=5 is well outside the range of the original data

​(b) Predict the ERA for

10

​wins, if it is meaningful. Select the correct choice below​ and, if​ necessary, fill in the answer box within your choice.

A) y=___ round to 2 decimal places if needed

B) it is not meaningful... same option as above for be only with 10 instead

C) it is not meaningful... same option as for c above except it says x=10

​(c) Predict the ERA for

twenty one

​wins, if it is meaningful. Select the correct choice below​ and, if​ necessary, fill in the answer box within your choice.

same options as above for a b and c except wit the 21

​(d) Predict the ERA for

fifteen

​wins, if it is meaningful. Select the correct choice below​ and, if​ necessary, fill in the answer box within your choice.

.18
Answer #1

I HOPE ITS HELPFUL TO YOU IF YOU HAVE ANY DOUBTS PLS COMMENTS BELOW..I WILL BE THERE TO HELP YOU...ALL THE BEST...

ANSWER:

Here i am taking this values

"Wins, x","Earned run average, y"
20,2.78
18,3.23
17,2.67
16,3.73
14,3.89
12,4.39
11,3.85
9,5.09

Here is the grafh

4 R2 = 0.8055 (Y) _ Linear 0 0 10 15 20 25

0.1925r +6.5195 y ー -

a) When x = 5

yーー0.1925 * 5 + 6.5195 5.557

b) When x = 10

0.1925 * 10 + 6.5195 = 4.5945 y.

c) When x = 21

y =-0.1925 * 21 6.5195 = 2.477

d) When x =15

y--0.1925*15 +6.5195-3.632

HOPE YOU UNDERSTAND...

PLS RATE THUMBS UP....ITS HELPS ME ALOT...

THANK YOU...!!